# GAN decoder on a quantum toric code for noise-robust quantum teleportation

### v2

@@ V Conclusion
In this work, we have employed a GAN algorithm to train and obtain a suited quantum toric code decoder. For toric code lattices of scale $d=3$ and $d=5$, the error correction success rates of this model reaches 99.739% and 99.895%, respectively, at an error rate $p=0.05$. Compared to the fidelity threshold of 0.1099 for the MWPM model, our model achieves a fidelity threshold of $P=0.2108$, demonstrating superior fidelity performance.
@@
Future research could further optimize GAN decoders with code distances of $d=7$ and beyond. Additionally, exploring various GAN architectures, such as Quantum Generative Adversarial Networks (QGANs) [49, 50, 51], CycleGAN [52], Self-Attention GAN (SAGAN) [53], may improve the efficiency and accuracy of decoders, thereby achieving higher fidelity in a broader range of quantum computing applications.
@@ Acknowledgements.
The authors acknowledge the Natural Science Foundation of Shandong Province of China (ZR2021ZD19). R.L.F. acknowledges support by MUR (Ministero dell’Università e della Ricerca) through the following projects: PNRR Project ICON-Q – Partenariato Esteso NQSTI – PE00000023 – Spoke 2 – CUP: J13C22000680006, PNRR Project QUANTIP – Partenariato Esteso NQSTI – PE00000023 – Spoke 9 – CUP: E63C22002180006, PNRR Project AQuSDIT – Partenariato Esteso SERICS – PE00000014 – Spoke 5 – H73C22000880001, PNRR Project PRISM – Partenariato Esteso RESTART – PE00000001 – Spoke 4 – CUP: C79J24000190004. Besides, Program of China Scholarship Council (Grant NO. 202306330060) are gratefully acknowledged.

### v3

@@ V Conclusion
In this work, we have employed a GAN algorithm to train and obtain a suited quantum toric code decoder. For toric code lattices of scale $d=3$ and $d=5$, the error correction success rates of this model reaches 99.739% and 99.895%, respectively, at an error rate $p=0.05$. Under the same conditions, our model demonstrates superior fidelity compared to the MWPM decoder. Moreover, it achieves a pseudo-threshold of $p=0.2108$, which is substantially higher than that of the MWPM decoder (0.1099).
@@
Regarding scalability, the proposed GAN-based decoder can, in principle, be extended to larger code distances by retraining or fine-tuning the model on appropriately generated datasets. The generator–discriminator structure enables parallel inference, which may help mitigate the increase in decoding latency as the system size grows. However, scaling to very large distances will inevitably increase the input dimensionality and model parameters, which could lead to higher training costs and memory consumption. To address this, future work may explore hierarchical GAN architectures or dimensionality reduction techniques to preserve decoding performance while maintaining computational efficiency. Additionally, exploring various GAN architectures, such as Quantum Generative Adversarial Networks (QGANs) [51, 52, 53], CycleGAN [54], Self-Attention GAN (SAGAN) [55], may improve the efficiency and accuracy of decoders, thereby achieving higher fidelity in a broader range of quantum computing applications.
@@ Acknowledgements.
The authors acknowledge the Natural Science Foundation of Shandong Province of China (ZR2021ZD19). R.L.F. acknowledges support by MUR (Ministero dell’Università e della Ricerca) through the following projects: PNRR Project ICON-Q – Partenariato Esteso NQSTI – PE00000023 – Spoke 2 – CUP: J13C22000680006, PNRR Project QUANTIP – Partenariato Esteso NQSTI – PE00000023 – Spoke 9 – CUP: E63C22002180006, PNRR Project AQuSDIT – Partenariato Esteso SERICS – PE00000014 – Spoke 5 – CUP: H73C22000880001, PNRR Project PRISM – Partenariato Esteso RESTART – PE00000001 – Spoke 4 – CUP: C79J24000190004. Besides, Program of China Scholarship Council (Grant NO. 202306330060) are gratefully acknowledged.